If the ratio of the ages of Kissi and Esinam is 3:5 and that of Esinam and Lariba is 3:5 and
the sum of the ages of all 3 is 147 years, what is the age difference between oldest the
youngest?
1
Expert's answer
2020-06-10T18:26:57-0400

Let the age of Kissi ,Esinam,Lariba are "x,y,z" respectively.

Then According to question,

"x:y=3:5 \ , \ y:z=3:5"

And "x+y+z=147"

Since, "x:y=3:5 \implies \frac{x}{y}=\frac{3}{5} \implies x=\frac{3y}{5}"

"y:z=3:5 \implies \frac{y}{z}=\frac{3}{5} \implies z=\frac{5y}{3}"


"\therefore \frac{3y}{5}+y+\frac{5y}{3}=147"


"\implies \frac{9y+15y+25y}{15}=147"


"\implies 49y=147×15"


"\implies y=\frac{147×15}{49}=3×15=45"


"\therefore x=\frac{3×45}{5}=27 \ and \ z=\frac{5×45}{3}=5×15=75"

The age difference between oldest and youngest "=75-27=48"